# Protocol as Poetry: Case Study on Pak’s Protocol Arts

### v1

@@ A.5. Lost Poets
Lost Poets launched with 65,536 Page NFTs at \$0.32 ETH each, selling out and raising roughly \$20 million in its initial sale. 1,024 Origin Poets were distributed (294 to top holders, 730 via daily random drops) . Phase II saw thousands of Pages burned: within the first week, over 50,000 Pages were converted into Poet NFTs, leaving near 15,000 Pages for the final phase. Each Poet NFT is programmatically one of 1,024 “families” and has 256 trait parameters , making each uniquely identifiable. As of the end of Phase III, about 15,000 Poets received names and poems (meaning that many Pages were sacrificed to finalize them), while a number of Poets remained “silent” because their owners chose not to burn pages. The most popular names taken include historical luminaries (“Shakespeare” was named within minutes of launch). Secondary market activity was robust: Origin Poets traded at a premium (some over 5 ETH each) due to their limited supply and role as progenitors. By project’s end, the community had collectively written tens of thousands of individual one-line poems, effectively creating one of the largest collaborative literary works in NFT form. Lost Poets won the 2022 NFT Award for Best Interactive Project, underscoring its impact on participatory art.
@@ A.6. Hate / Invisible Mechanisms (2021)
Figure 8. Comment in the source code of smart contract of Hate NFT, Invisible Mechanisms (2021)
@@
Invisible Mechanism are intertwined artworks that together form a provocative commentary on control, ownership, and the relationship between creator and audience (even the antagonistic part of the audience). In late 2021, Pak executed Hate as a conceptual performance: he offered special NFTs titled “Hate” to thirty of his harshest critics on social media – effectively gifting tokens to those who had publicly disparaged him (Pak, 2021). The twist was that these Hate NFTs were programmed to be immutable and non-transferable by the recipients, symbolizing the idea that their negativity had earned them a token they could not get rid of. This gesture served as both a prank and a statement, turning the act of hating Pak into an involuntary participation in his art. Underpinning Hate was Pak’s new smart contract prototype called Move: an “invisible mechanism” that grants the creator the power to transfer any token from any wallet within that contract to another address at will, irrespective of holder consent. Move embodies the mantra “CREATION OBEYS ITS CREATOR”, flipping the usual decentralization narrative to assert total artistic control over distribution. Together, Hate (the scenario) and Move (the technology) form a conceptual piece that asks pointed questions: What if an artist could reclaim authority over an artwork after it’s been distributed? Can an artwork be a form of dialogue – or retaliation – between artist and critic? By engaging directly with haters and by unveiling a contract that subverts NFT norms, Pak’s concept challenged the community to reflect on power dynamics and the presumed inviolability of ownership in the crypto space.
@@
The Hate NFTs were engineered using the Move contract architecture, which fundamentally alters standard ERC-721 behavior (202, 2021a). Normally, once an NFT is in a wallet, only the owner can transfer it; Pak’s custom contract introduced admin functions that allowed him (as the contract owner) to lock tokens and later move them at will . Upon minting the 30 Hate tokens, they were immediately “locked” to their respective recipient addresses – hence any transfer or sale initiated by those holders would automatically fail, as the contract overrode such actions . The NFTs had identical metadata (depicted simply by a heart symbol) , emphasizing that their differentiating feature was the wallet they resided in, not visual content. After letting the scenario play out for about 24 hours, Pak utilized Move to demonstrate its power: he, as admin, transferred all Hate tokens out of the haters’ wallets back into his own, even updating their metadata before doing so . This action conclusively illustrated the dynamic – the artist could literally reclaim the tokens at will, something impossible under standard token contracts. The Move contract thus created a dynamic where token ownership was provisional, entirely subject to the creator’s whims. Such a mechanism is invisible until exercised, earning its name, and it turned the typical trust model of NFTs on its head: here, the code enforced a hierarchy where the art always ultimately “belonged” to its maker.
@@
Interestingly, Hate inverted typical participation – it was the artist who “participated” in the lives of certain audience members by depositing NFTs into their wallets uninvited. Pak solicited volunteers by asking his followers to prove they had posted genuine hate toward him (via old tweets) . The “winners” of this dubious honor were then airdropped the Hate NFTs; about 30 individuals received one . Their engagement was largely one of surprise and frustration: many attempted to trade or remove the tokens, only to discover every transfer attempt was automatically rejected by the contract . In essence, their interaction was to experience powerlessness – an unusual, perhaps uncomfortable form of audience participation where the usual agency of a collector was stripped away. As for Move, the wider community’s participation came through observation and discussion rather than direct use (since Move is a contract design, not a direct-to-consumer app). Pak publicly shared the Move contract and its implications for all to inspect , inviting creators and collectors to contemplate this new paradigm. In doing so, the community collectively engaged in a discourse about consent and control in NFTs. In summary, Hate/Move turned the tables on engagement: instead of people actively opting into the art, the art opted into their lives, thereby making the audience’s reactions and discussions (especially the haters’ public bemusement) an integral part of the piece.
@@
By its very nature, Hate/Move was an exercise in maximizing artist control. Pak retained absolute authority over the Hate tokens at all times – he was the only one who could mint them, the only one who could move or burn them, and even the only one who could change their content . This project was a deliberate outlier compared to his other works: instead of surrendering control to autonomous rules or the community, he crafted a scenario where his hand remained firmly on the lever from start to finish. In fact, the project can be seen as a manifesto of control: Move was Pak asserting that as an artist-programmer, he could imbue a token with a piece of his will, overriding the default freedoms typically given to collectors. After the initial performative phase, Pak ended the experiment on his terms by retrieving the tokens and effectively erasing them from the participants’ wallets . In doing so, he underlined the central statement: these creations were never truly “theirs” – they were on loan at the creator’s mercy. Such extreme control is generally anathema to the decentralization ethos, which is precisely why Hate/Move is so incisive; it used exaggerated artist control as an artistic device itself.
Collective Emerging Behavior
The immediate participants – the recipients of Hate – reacted with a mix of confusion, amusement, and annoyance. Some joked about selling their entire wallet (since they couldn’t sell the token itself), highlighting the absurd lengths one would need to go to bypass the restriction . This small group of 30 became inadvertent performers in Pak’s drama, their attempts and public commentary forming a crucial part of the narrative that others followed. Meanwhile, the broader NFT community watched the spectacle with fascination. On social media and blogs, a flurry of debate ensued about the ethics and implications of Move: some collectors expressed relief that such a mechanism was not widespread, while some creators were intrigued by the power it demonstrated. The collective behavior here was largely discursive – Hate/Move spurred conversations about the nature of ownership and the extent of smart contract programmability. People questioned how this contrasts with the spirit of self-sovereignty that NFTs usually promise, and whether it was “fair” or simply a clever conceptual stunt. In the end, once Pak reclaimed the Hate NFTs, the haters were free of their burdens, perhaps a bit wiser to the power of code in digital art. The community that observed gained a nuanced understanding: emerging from it was a collective acknowledgement that decentralization in art is not absolute, and that the artist’s intent coded into a contract can drastically shape user experience. This dialogue and reflection were exactly the emergent cultural outcomes Pak likely sought, elevating the work from a mere prank to a significant case study in the NFT space.
@@
In retrospect, Hate/Move reads like a sharp, if mischievous, parable within the NFT art narrative. It poetically frames the tension between artist and audience: the Hate tokens were as much a mirror to the recipients (reflecting their hostility back as an unusable “gift”) as they were a canvas for Pak’s statement about artistic sovereignty. There’s dark humor and irony in the idea that hatred towards an artist could be alchemized into an artwork that essentially traps that hate – a modern digital twist on holding up a mirror to one’s critics. The Move mechanism adds an extra layer of meaning: it challenged the community to realize that the liberties they take for granted in decentralized art can be tweaked or overturned by a clever creator. This provocation was unsettling to some and thrilling to others, serving as what one might call a necessary thought experiment in a hype-driven field. By co-opting his detractors into unwilling collaborators, Pak turned negativity into a generative component of his practice, raising questions about consent, power, and the very definition of ownership. In the broader context of digital art discourse, Hate/Move is significant for revealing that the blockchain is not inherently liberating – its effect depends on how it’s used. Pak’s broader impact here is to remind both creators and collectors that the medium’s rules can themselves be the artwork. Hate/Move ultimately stands as a conceptual punchline with a serious core: an artwork about authority, delivered in the medium of authority itself.
@@
In November 2021, Pak airdropped 30 one-of-a-kind Hate NFTs (depicted as a simple black square with the word “HATE”) to selected critics . Each had an embedded “move” mechanism allowing Pak to transfer them. The contract (address 0xMove…) was written by Manifold Studio and made public on Etherscan . During the first week, at least 2 instances of Pak using the move function were observed, and 0 transfers by holders (since none could occur). The stunt quickly spread awareness; tweets about “Pak’s Hate” garnered thousands of impressions, and articles in crypto media (e.g. CryptoTimes and others) explained the phenomenon . The term “Invisible Mechanism” was coined by Pak in a tweet announcing the project’s true purpose . Ultimately, no monetary exchange took place for these NFTs on the market (until perhaps after Pak later unlocked them, if he did). Instead, their value was purely conceptual. Pak’s “Move” contract was later referenced in discussions on NFT standards, making these 30 tokens and their story a small but significant footnote in NFT history about creative smart contract design.
@@ A.7. Merge
Figure 9. Token mechanism shown on the Merge (2021) launch website.
Figure 10. Third-party interactive data visualization of Merge (2021) at https://mass.show/
@@
Merge (December 2021) is a landmark NFT work wherein Pak explored themes of unity, rarity, and collective participation by creating an artwork that exists in fragments yet has the potential to become a single whole. Billed as the first artwork that collectors could collectively assemble, Merge was sold not as discrete editions but as “mass units” – small identical tokens of mass that buyers could acquire in any amount during a 48-hour sale. Each buyer’s mass units automatically fused into one NFT (a single mass) in their wallet, whose size (visualized by a growing circular form) was proportional to the amount of mass purchased. The concept plays on the idea of merging: if two Merge NFTs ever end up in the same wallet, they combine into one larger mass, reducing the total number of tokens in circulation. In effect, Pak conceived Merge as a dynamic, self-collapsing collection – one that challenged the traditional notion of an edition size by making the supply theoretically shrinkable over time. Thematically, Merge invokes a sense of digital unity: all collectors hold a piece of a conceptually singular artwork, and if one entity were to gather every piece, the title’s promise would be realized as a single merged entity. Through this, Pak commented on the interplay between collaboration and competition, inviting the question of whether collectors would consolidate or guard their individual pieces, and what ultimate form the artwork might take.
@@
The Merge contract innovated on the token model by introducing an additive property to NFTs. Technically, each mass unit was an ERC-1155 token minted during the sale; at the close of the sale, the contract “compressed” a buyer’s multiple units into a single ERC-721 NFT that recorded the total mass count as an attribute. The NFT’s visual appearance (a circle of a certain size and perhaps color) was programmatically determined by the mass count, making the artwork generative and data-driven. The merging logic was enforced at the contract level: it utilized a transfer hook that, upon detecting a Merge token arriving in a wallet that already held one, would cancel the separate existence of the incoming token and increment the mass of the token already in that wallet . This mechanism guaranteed that no wallet could ever hold more than one Merge NFT – a radical departure from standard NFTs. Over time, if collectors consolidated holdings or accidentally merged by buying with a pre-existing mass in their wallet, the total token count of the project would decrease from the initial 28,983 supply. Unlike most art, where the edition size is fixed, Merge had a fluid supply that could contract, theoretically even to 1. This dynamic introduced emergent phenomena: for example, the rarity of certain visual variants depended on how people managed their tokens, not solely on predetermined traits. By embedding these rules, Pak effectively encoded a set of interactions and potential outcomes (even conflict) into the artwork’s DNA, making Merge as much a social experiment as a digital sculpture.
@@
Participation in Merge was straightforward yet unprecedented in scale: 28,983 collectors took part in the open sale, collectively purchasing 312,686 mass units over the two-day period . Instead of competing for limited editions, buyers were cooperating in a sense – everyone was guaranteed to receive their own NFT mass, and many aimed to accumulate as much mass as they could. The interactive twist emerged after the sale: whenever a Merge NFT was transferred on the secondary market, the contract would check the recipient’s wallet. If the recipient already had a Merge NFT, the incoming mass would merge with the existing one, destroying one token and increasing the other’s mass . This meant that collectors had to strategize; some chose to split their holdings across multiple wallets to keep separate smaller masses (especially if those had unique visual traits like a different color), while others intentionally merged to build a singular, more “massive” mass. There was even a gamified angle of aggression and defense: theoretically, someone could send a large mass to another collector’s address (without permission) to forcibly merge and absorb a smaller mass, as a way to sabotage its rarity or uniqueness . Such scenarios made holding a Merge NFT a participatory experience beyond the initial purchase: the community was actively discussing merger tactics, bragging about the size of their mass, and contemplating alliances or friendly competitions. In essence, Merge turned collecting into a collective game that extended into every transaction post-drop.
@@
After Merge was launched, Pak’s role was largely hands-off; the rules in the smart contract dictated the fate of the tokens. He did not exert control over who could buy or how much (beyond setting the sale’s time frame and pricing increments), nor could he intervene in the merging process once tokens were in circulation – that process was entirely automatic and irrevocable. However, his control was asserted in the careful design: he set the initial conditions that governed everything, from the pricing mechanism (which increased the price per mass unit in stages as certain sales milestones were hit) to the visual algorithm that represented mass. By relinquishing active control and entrusting the artwork to self-execution, Pak reinforced Merge’s stance as a decentralized piece governed by protocol. In fact, the only way Pak influenced the outcome was indirectly, through his initial decision not to cap the edition: this allowed the community’s demand to define the scale of the piece (and ultimately break the record for the largest-ever art sale by a living artist). In the life of Merge post-sale, Pak stepped back completely – any further evolution (like tokens merging into fewer tokens) was purely at the discretion of the holders and the mechanics he’d put in place. This minimal ongoing control is consistent with Pak’s tendency to let his protocol-based works run their course autonomously, in stark contrast to the interventionist stance of Hate/Move.
Collective Emerging Behavior

### v2

@@ A.5. Lost Poets
Lost Poets launched with 65,536 Page NFTs at 0.32 ETH each, selling out and raising roughly \$70 million in its initial sale. 1,024 Origin Poets were distributed (277 to top 100 collectors, 730 via daily random drops, 17 reserved). Phase II saw thousands of Pages burned: within the first week, over 50,000 Pages were converted into Poet NFTs, leaving near 15,000 Pages for the final phase. Each Poet NFT is programmatically one of 1,024 “families” and has 256 trait parameters, making each uniquely identifiable. As of the end of Phase III, about 15,000 Poets received names and poems (meaning that many Pages were sacrificed to finalize them), while a number of Poets remained “silent” because their owners chose not to burn pages. The most popular names taken include historical luminaries (“Shakespeare” was named within minutes of launch). Secondary market activity was robust: Origin Poets traded at a premium (some over 5 ETH each) due to their limited supply and role as progenitors. By project’s end, the community had collectively written tens of thousands of individual one-line poems, effectively creating one of the largest collaborative literary works in NFT form.
Annotation
Lost Poets is the richest case for C3 (distributed agency) and C4 (temporal dynamism): its multi-act structure unfolds over a full year in a genuinely performative and crowdsourced process, and collectors become genuine co-authors by naming their Poets and choosing whether to transform, hold, or burn their tokens. The temporal dynamic is essential: each phase—from the initial Page distribution through the Reveal, the Explorer phase, and the final Twist—constitutes a distinct participatory epoch, and the work’s meaning accrues through the cumulative decisions of thousands of participants across these stages. It also deeply exemplifies C6 (poetic message embedding), with the Borges-inspired narrative framing the naming ritual as a meditation on language and memory. C7 (composability) is significant: the project bridges into the \$ASH ecosystem, utilizing the burn mechanism introduced in burn.art—participants can burn Poets for \$ASH, and \$ASH holders received airdropped Pages, creating a feedback loop between projects. What makes Lost Poets distinctive is the depth of its branching decision space—at each stage, every participant faces a choice to transfer, hold, or burn, and each holder’s path through the Acts yields a different outcome, producing an emergent diversity that exceeds what any single-mechanism work (like X or Merge) can generate. The project demonstrates that everyone can participate in the unfolding of protocol art, and that the temporally extended, multi-stage structure is essential for cultivating the kind of sustained collective authorship that transforms a smart contract into a living literary work.
@@ A.6. Hate / Invisible Mechanisms (2021)
Refer to caption Large ASCII-art text rendered in monospace characters on a dark background spelling out “CREATION OBEYS ITS CREATOR” in three lines. Each letter is composed of smaller text characters that appear to be fragments of smart contract source code, visually fusing the philosophical statement with the underlying program code.
Figure 8. “CREATION OBEYS ITS CREATOR”—a statement embedded as ASCII art in the source code of the Hate NFT smart contract, part of Pak’s Invisible Mechanisms (2021). The phrase encapsulates the Move protocol’s radical inversion of NFT ownership: the creator retains administrative power to lock, transfer, or reclaim any token at will.
@@
Invisible Mechanism are intertwined artworks that together form a provocative commentary on control, ownership, and the relationship between creator and audience (even the antagonistic part of the audience). In late 2021, Pak executed Hate as a conceptual performance: Pak offered special NFTs titled “Hate” to thirty of Pak’s harshest critics on social media—effectively gifting tokens to those who had publicly disparaged Pak (The Crypto Times, 2021b). The twist was that these Hate NFTs were programmed to be immutable and non-transferable by the recipients, symbolizing the idea that their negativity had earned them a token they could not get rid of. This gesture served as both a prank and a statement, turning the act of hating Pak into an involuntary participation in Pak’s art. Underpinning Hate was Pak’s new smart contract prototype called Move: an “invisible mechanism” that grants the creator the power to transfer any token from any wallet within that contract to another address at will, irrespective of holder consent. Move embodies the mantra “CREATION OBEYS ITS CREATOR”, flipping the usual decentralization narrative to assert total artistic control over distribution. Together, Hate (the scenario) and Move (the technology) form a conceptual piece that asks pointed questions: What if an artist could reclaim authority over an artwork after it’s been distributed? Can an artwork be a form of dialogue—or retaliation—between artist and critic? By engaging directly with haters and by unveiling a contract that subverts NFT norms, Pak’s concept challenged the community to reflect on power dynamics and the presumed inviolability of ownership in the crypto space.
@@
The Hate NFTs were engineered using the Move contract architecture, which fundamentally alters standard ERC-721 behavior (The Crypto Times, 2021a). Normally, once an NFT is in a wallet, only the owner can transfer it; Pak’s custom contract introduced admin functions that allowed Pak (as the contract owner) to lock tokens and later move them at will. Upon minting the 30 Hate tokens, they were immediately “locked” to their respective recipient addresses—hence any transfer or sale initiated by those holders would automatically fail, as the contract overrode such actions. The NFTs had identical metadata (depicted simply by a heart symbol), emphasizing that their differentiating feature was the wallet they resided in, not visual content. After letting the scenario play out for about 24 hours, Pak utilized Move to demonstrate its power: Pak, as admin, transferred all Hate tokens out of the haters’ wallets back into Pak’s own, even updating their metadata before doing so. This action conclusively illustrated the dynamic—the artist could literally reclaim the tokens at will, something impossible under standard token contracts. The Move contract thus created a dynamic where token ownership was provisional, entirely subject to the creator’s whims. Such a mechanism is invisible until exercised, earning its name, and it turned the typical trust model of NFTs on its head: here, the code enforced a hierarchy where the art always ultimately “belonged” to its maker.
@@
Hate inverted typical participation—it was the artist who “participated” in the lives of certain audience members by depositing NFTs into their wallets uninvited. Pak solicited volunteers by asking Pak’s followers to prove they had posted genuine hate toward Pak (via old tweets). The “winners” of this dubious honor were then airdropped the Hate NFTs; about 30 individuals received one. Their engagement was largely one of surprise and frustration: many attempted to trade or remove the tokens, only to discover every transfer attempt was automatically rejected by the contract. In essence, their interaction was to experience powerlessness—an unusual, perhaps uncomfortable form of audience participation where the usual agency of a collector was stripped away. As for Move, the wider community’s participation came through observation and discussion rather than direct use (since Move is a contract design, not a direct-to-consumer app). Pak publicly shared the Move contract and its implications for all to inspect, inviting creators and collectors to contemplate this new paradigm. In doing so, the community collectively engaged in a discourse about consent and control in NFTs. In summary, Hate/Move turned the tables on engagement: instead of people actively opting into the art, the art opted into their lives, thereby making the audience’s reactions and discussions (especially the haters’ public bemusement) an integral part of the piece.
@@
By its very nature, Hate/Move was an exercise in maximizing artist control. Pak retained absolute authority over the Hate tokens at all times—Pak was the only one who could mint them, the only one who could move or burn them, and even the only one who could change their content. This project was a deliberate outlier compared to Pak’s other works: instead of surrendering control to autonomous rules or the community, Pak crafted a scenario where Pak’s hand remained firmly on the lever from start to finish. In fact, the project can be seen as a manifesto of control: Move was Pak asserting that as an artist-programmer, Pak could imbue a token with a piece of Pak’s will, overriding the default freedoms typically given to collectors. After the initial performative phase, Pak ended the experiment on Pak’s own terms by retrieving the tokens and effectively erasing them from the participants’ wallets. In doing so, Pak underlined the central statement: these creations were never truly “theirs”—they were on loan at the creator’s mercy. Such extreme control is generally anathema to the decentralization ethos, which is precisely why Hate/Move is so incisive; it used exaggerated artist control as an artistic device itself.
Collective Emergent Behavior
The immediate participants—the recipients of Hate—reacted with a mix of confusion, amusement, and annoyance. Some joked about selling their entire wallet (since they couldn’t sell the token itself), highlighting the absurd lengths one would need to go to bypass the restriction. This small group of 30 became inadvertent performers in Pak’s drama, their attempts and public commentary forming a crucial part of the narrative that others followed. Meanwhile, the broader NFT community watched the spectacle with fascination. On social media and blogs, a flurry of debate ensued about the ethics and implications of Move: some collectors expressed relief that such a mechanism was not widespread, while some creators were intrigued by the power it demonstrated. The collective behavior here was largely discursive—Hate/Move spurred conversations about the nature of ownership and the extent of smart contract programmability. People questioned how this contrasts with the spirit of self-sovereignty that NFTs usually promise, and whether it was “fair” or simply a clever conceptual stunt. In the end, once Pak reclaimed the Hate NFTs, the haters were free of their burdens, perhaps a bit wiser to the power of code in digital art. The community that observed gained a nuanced understanding: emerging from it was a collective acknowledgement that decentralization in art is not absolute, and that the artist’s intent coded into a contract can drastically shape user experience. This dialogue and reflection were exactly the emergent cultural outcomes Pak likely sought, elevating the work from a mere prank to a significant case study in the NFT space.
@@
In retrospect, Hate/Move reads like a sharp, if mischievous, parable within the NFT art narrative. It poetically frames the tension between artist and audience: the Hate tokens were as much a mirror to the recipients (reflecting their hostility back as an unusable “gift”) as they were a canvas for Pak’s statement about artistic sovereignty. There’s dark humor and irony in the idea that hatred towards an artist could be alchemized into an artwork that essentially traps that hate—a modern digital twist on holding up a mirror to one’s critics. The Move mechanism adds an extra layer of meaning: it challenged the community to realize that the liberties they take for granted in decentralized art can be tweaked or overturned by a clever creator. This provocation was unsettling to some and thrilling to others, serving as what one might call a necessary thought experiment in a hype-driven field. By co-opting Pak’s detractors into unwilling collaborators, Pak turned negativity into a generative component of Pak’s practice, raising questions about consent, power, and the very definition of ownership. In the broader context of digital art discourse, Hate/Move is significant for revealing that the blockchain is not inherently liberating—its effect depends on how it’s used. Pak’s broader impact here is to remind both creators and collectors that the medium’s rules can themselves be the artwork. Hate/Move ultimately stands as a conceptual punchline with a serious core: an artwork about authority, delivered in the medium of authority itself.
@@
In November 2021, Pak airdropped 30 one-of-a-kind Hate NFTs (depicted as a simple heart symbol) to selected critics. Each had an embedded “move” mechanism allowing Pak to transfer them. The contract (address 0xMove…) was written by Manifold Studio and made public on Etherscan. During the first week, at least 2 instances of Pak using the move function were observed, and 0 transfers by holders (since none could occur). The stunt quickly spread awareness; tweets about “Pak’s Hate” garnered thousands of impressions, and articles in crypto media (e.g., CryptoTimes and others) explained the phenomenon. The term “Invisible Mechanism” was coined by Pak in a tweet announcing the project’s true purpose. Ultimately, no monetary exchange took place for these NFTs on the market (until perhaps after Pak later unlocked them, if Pak did). Instead, their value was purely conceptual. Pak’s “Move” contract was later referenced in discussions on NFT standards, making these 30 tokens and their story a small but significant footnote in NFT history about creative smart contract design.
Annotation
Invisible Mechanisms is an outlier that inverts nearly every pattern in the portfolio. Where other works distribute control (C2), this one concentrates it absolutely: the move() function grants the artist unilateral power to seize or relocate any token. Where other works invite willing participation (C3), here the recipients were enlisted without full knowledge of the token’s constraints. The work’s true innovation, however, lies in its creation of a non-transferable asset—a token that, once deposited in a wallet, cannot be moved by its holder. This is a radical contribution to C1 (system-centric composition): the contract’s permissions structure is the artistic statement, and the meaning of the work resides entirely in what the protocol forbids rather than what it enables. The non-transferability mechanism directly anticipates the “soulbound token” concept later formalized by Weyl, Ohlhaver, and Buterin (Weyl et al., 2022), making Hate a historically significant precedent in which an artistic experiment prefigured a major theoretical development in blockchain design. The work’s primary contribution is to C6 (poetic message embedding): the non-transferable “Hate” token literalizes the idea that negative attention, once given, cannot be taken back—a pointed commentary on power asymmetry encoded directly in the smart contract. Furthermore, the mechanism pioneered here did not remain isolated: the non-transferability logic later inspired the soul-bound design deployed in Censored, where message tokens were locked until Assange’s release, demonstrating that even Pak’s most provocative experiments serve as compositional building blocks for subsequent works.
@@ A.7. Merge
Refer to caption A screenshot from the Merge launch website. At top, a 3D rendering of two white blob-like spheres pressing together, illustrating the merging concept. Below, the heading “Token mechanism” followed by explanatory text: “Merge Collection has a built-in scarcity mechanism to ensure token supply decreases over time. Every Merge token transfer merges it with the token in the recipient wallet, adding up the Mass value and resulting in a single token.” At the bottom, the formula m(100) plus m(20) equals m(120) is displayed.
Figure 9. The token mechanism explained on the Merge (2021) launch website. When two tokens enter the same wallet, the smart contract automatically combines them into one with summed mass—a built-in deflationary protocol that ensures the total supply can only decrease over time.
A screenshot of the Merge collection on OpenSea showing approximately twenty NFTs displayed as colored circles of varying sizes on colored backgrounds. Circles range from small to very large, colored white, yellow, red, and blue against blue, purple, or black backgrounds. The header shows collection stats: 27,307 items, floor price 0.0326 ETH, total volume 2,975 ETH. A sidebar displays filtering options for Tier (1 through 4), Merges count, Mass, Class, and Alpha traits.
Figure 10. Merge on OpenSea. Each NFT is visualized as a circle whose size corresponds to its accumulated mass. Color tiers (white, yellow, red, blue) indicate mass thresholds, making the relative “weight” of each collector’s holdings immediately legible across the marketplace.
A third-party data visualization resembling a galaxy or solar system. Thousands of small white dots representing individual Merge tokens are arranged in a radial pattern around a cluster of larger colored spheres (white, red, yellow, blue) at the center representing the highest-mass tokens. A timeline bar along the bottom shows temporal data. The overall effect evokes a cosmic distribution of mass.
Figure 11. Third-party interactive visualization of all Merge tokens at https://mass.show/. Each dot represents one token, sized proportionally to its mass. The galaxy-like distribution reveals the power-law concentration of mass among a few dominant holders surrounded by thousands of smaller participants.
@@
Merge (December 2021) is a landmark NFT work wherein Pak explored themes of unity, rarity, and collective participation by creating an artwork that exists in fragments yet has the potential to become a single whole. Billed as the first artwork that collectors could collectively assemble, Merge was sold not as discrete editions but as “mass units”—small identical tokens of mass that buyers could acquire in any amount during a 48-hour sale. Each buyer’s mass units automatically fused into one NFT (a single mass) in their wallet, whose size (visualized by a growing circular form) was proportional to the amount of mass purchased. The concept plays on the idea of merging: if two Merge NFTs ever end up in the same wallet, they combine into one larger mass, reducing the total number of tokens in circulation. In effect, Pak conceived Merge as a dynamic, self-collapsing collection—one that challenged the traditional notion of an edition size by making the supply theoretically shrinkable over time. Thematically, Merge invokes a sense of digital unity: all collectors hold a piece of a conceptually singular artwork, and if one entity were to gather every piece, the title’s promise would be realized as a single merged entity. Through this, Pak commented on the interplay between collaboration and competition, inviting the question of whether collectors would consolidate or guard their individual pieces, and what ultimate form the artwork might take.
@@
The Merge contract innovated on the token model by introducing an additive property to NFTs. Technically, each mass unit was an ERC-1155 token minted during the sale; at the close of the sale, the contract “compressed” a buyer’s multiple units into a single ERC-721 NFT that recorded the total mass count as an attribute. The NFT’s visual appearance (a circle of a certain size and perhaps color) was programmatically determined by the mass count, making the artwork generative and data-driven. The merging logic was enforced at the contract level: it utilized a transfer hook that, upon detecting a Merge token arriving in a wallet that already held one, would cancel the separate existence of the incoming token and increment the mass of the token already in that wallet. This mechanism guaranteed that no wallet could ever hold more than one Merge NFT—a radical departure from standard NFTs. Over time, if collectors consolidated holdings or accidentally merged by buying with a pre-existing mass in their wallet, the total token count of the project would decrease from the initial 28,983 supply. Unlike most art, where the edition size is fixed, Merge had a fluid supply that could contract, theoretically even to 1. This dynamic introduced emergent phenomena: for example, the rarity of certain visual variants depended on how people managed their tokens, not solely on predetermined traits. By embedding these rules, Pak effectively encoded a set of interactions and potential outcomes (even conflict) into the artwork’s DNA, making Merge as much a social experiment as a digital sculpture.
@@
Participation in Merge was straightforward yet unprecedented in scale: 28,983 collectors took part in the open sale, collectively purchasing 312,686 mass units over the two-day period. Instead of competing for limited editions, buyers were cooperating in a sense—everyone was guaranteed to receive their own NFT mass, and many aimed to accumulate as much mass as they could. The interactive twist emerged after the sale: whenever a Merge NFT was transferred on the secondary market, the contract would check the recipient’s wallet. If the recipient already had a Merge NFT, the incoming mass would merge with the existing one, destroying one token and increasing the other’s mass. This meant that collectors had to strategize; some chose to split their holdings across multiple wallets to keep separate smaller masses (especially if those had unique visual traits like a different color), while others intentionally merged to build a singular, more “massive” mass. There was even a gamified angle of aggression and defense: theoretically, someone could send a large mass to another collector’s address (without permission) to forcibly merge and absorb a smaller mass, as a way to sabotage its rarity or uniqueness. Such scenarios made holding a Merge NFT a participatory experience beyond the initial purchase: the community was actively discussing merger tactics, bragging about the size of their mass, and contemplating alliances or friendly competitions. In essence, Merge turned collecting into a collective game that extended into every transaction post-drop.
@@
After Merge was launched, Pak’s role was largely hands-off; the rules in the smart contract dictated the fate of the tokens. Pak did not exert control over who could buy or how much (beyond setting the sale’s time frame and pricing increments), nor could Pak intervene in the merging process once tokens were in circulation—that process was entirely automatic and irrevocable. However, Pak’s control was asserted in the careful design: Pak set the initial conditions that governed everything, from the pricing mechanism (which increased the price per mass unit in stages as certain sales milestones were hit) to the visual algorithm that represented mass. By relinquishing active control and entrusting the artwork to self-execution, Pak reinforced Merge’s stance as a decentralized piece governed by protocol. In fact, the only way Pak influenced the outcome was indirectly, through Pak’s initial decision not to cap the edition: this allowed the community’s demand to define the scale of the piece (and ultimately break the record for the largest-ever art sale by a living artist). In the life of Merge post-sale, Pak stepped back completely—any further evolution (like tokens merging into fewer tokens) was purely at the discretion of the holders and the mechanics Pak had put in place. This minimal ongoing control is consistent with Pak’s tendency to let Pak’s protocol-based works run their course autonomously, in stark contrast to the interventionist stance of Hate/Move.
Collective Emergent Behavior